According to the ideal gas law, a 0.9188 mol sample of xenon gas in a 1.004 L container at 268.5 K should exert a pressure of 20.16 atm. What is the percent difference between the pressure calculated using the van der Waals' equation and the ideal pressure? For Xe gas, a = 4.194 L2atm/mol2 and b = 5.105×10-2 L/mol.  
Answer #1

n = 0.9188mol

V = 1.004L

T   = 268.5K

pressure exert from ideal gas = 20.16atm

(P + an^2/V^2)(V-nb) = nRT

(P + 4.194*(0.9188)^2/(1.004)^2)(1.004-0.9188*5.105*10^-2) = 0.9188*0.0821*268.5

(P + 3.5123)(0.95709526) = 20.25388938

P   = 17.65atm

% difference = /P ideal - P wad/*100/P ideal

                     = (20.16-17.65)*100/20.16

                     = 251/20.16   = 12.45% >>>>answer